# Sonnet: Spectral Operator Neural Network for Multivariable Time Series Forecasting

### v1

@@ E.2 Supplementary experiment results with sMAPE score
Table S9: Performance (average sMAPE across the $4$ test seasons) of iTransformer, Samformer, and PatchTST on the ILI forecasting tasks (ILI-ENG/US2/US9) with different modifications to the naïve attention mechanism. ‘$\neg$ Attn’ denotes the removal of the residual attention module, and FNet / FED / VDAB refer to using the attention modules proposed in FNet [43], FEDformer [62], and DeformTime [43], respectively. This table supplements Table 2 that reports MAEs. Best results are bolded and second best are underlined.
Attention $H=7$ days $H=14$ days $H=21$ days $H=28$ days
ENG US2 US9 ENG US2 US9 ENG US2 US9 ENG US2 US9
iTransformer 26.38 23.24 18.57 36.67 28.17 22.44 48.93 30.03 24.11 55.35 36.75 31.05
FNet 30.16 26.99 22.27 41.28 31.12 26.29 57.16 36.48 29.69 64.26 41.06 35.00
FED 40.21 32.58 26.25 48.37 37.86 30.68 67.76 50.94 40.97 73.86 54.82 44.33
VDAB 27.46 24.55 19.21 38.10 30.39 23.29 48.19 32.37 23.07 54.16 35.39 26.97
SamFormer 28.31 24.21 19.39 36.63 30.16 24.50 54.41 31.42 24.22 60.78 36.8 27.95
FNet 30.01 27.94 21.95 40.33 33.06 27.33 53.16 36.21 31.30 61.64 42.28 34.68
FED 40.18 32.35 26.34 48.91 37.65 30.84 67.79 49.99 41.26 73.17 53.91 44.85
VDAB 25.98 21.88 17.05 36.84 27.48 20.95 46.52 30.78 22.29 57.09 33.49 25.63
PatchTST 27.61 24.51 19.34 37.76 30.11 24.09 51.11 36.70 29.40 59.60 42.61 33.35
FNet 31.93 28.81 22.14 44.15 33.77 25.64 58.17 38.11 26.46 60.51 41.48 29.92
FED 40.21 32.59 26.25 50.26 37.88 30.76 67.85 50.89 41.16 74.08 54.61 44.42
VDAB 26.88 22.99 17.99 38.08 30.99 20.64 50.49 32.45 23.90 60.16 34.89 25.61
@@
E.3 Sensitivity to random seed initialisation
Table S10: Sensitivity to different seed initialisations ($5$ seeds) for Sonnet using the ILI-ENG data set across all forecasting horizons ($H$). $\mu$ and $\sigma$ denote the mean and standard deviation of the $5$ obtained MAEs per test period. The results in the main paper were obtained for seed ‘42’.
@@ E.3 Sensitivity to random seed initialisation
$\mu\leavevmode\nobreak\ (\sigma)$ 1.60 (0.07) 1.46 (0.14) 2.92 (0.07) 2.01 (0.20) 2.00 (0.07)
$H$ seed 2015/16 2016/17 2017/18 2018/19 Average
@@
$\mu\leavevmode\nobreak\ (\sigma)$ 2.23 (0.15) 1.69 (0.08) 4.40 (0.15) 2.88 (0.10) 2.80 (0.06)
Results that assess the sensitivity of Sonnet to different random seed initialisations are enumerated in Table S10. We have examined $5$ different seeds (‘10’, ‘42’, ‘111’, ‘$1$’, ‘1234’) while conducting experiments on the ILI-ENG data set for all test periods and forecasting horizons. The obtained results highlight that the proposed model provides consistent results across all regions and forecasting horizons. Results do not deviate significantly from the ones presented in the main paper (Table 1) which are obtained using seed ‘42’. Standard deviations in the averaged MAEs (all test periods) over the $5$ seeds for each forecasting horizon do not exceed $0.07$ which in real terms (using the units of the actual data) represents $7$ people in a population of $10$ million. This highlights the robustness of the mean estimate.
E.4 Analysis of model performance for tasks with strong seasonality
Table S11: Effect of different $\alpha$ values for Sonnet while conducting experiments with the ELEC data set. $r$ and $\epsilon\%$ denote the correlation and sMAPE, respectively. “Naïve” in the last row denotes a seasonal persistence model. Best results are bolded and second best are underlined.
@@ E.4 Analysis of model performance for tasks with strong seasonality
Prior work has shown that modelling inter-variable dependencies in time series problems may occasionally cause overfitting [34]. While our results for the ILI and WEA forecasting tasks do not corroborate this (section 4.2), we conduct further experiments on a different task (ELEC) that displays a different pattern of performance across most models. In these experiments, we explore different values for the parameter $\alpha$ in Sonnet. $\alpha$ controls the relative contribution of exogenous variables in the learned representations (section 3.1). For $\alpha=0$, Sonnet only uses historical values of the target (endogenous) variable. We use the ELEC data set for evaluation because for this specific task, PatchTST—that does not capture inter-variable dependencies—outperforms some models that do (e.g. ModernTCN and Crossformer). In addition, $7$ out of $9$ evaluated models cannot surpass the accuracy of the seasonal persistence model (both in terms of MAE and sMAPE), indicating that for ELEC, seasonal patterns are consistent over time. Sonnet’s hyperparameters are re-tuned (see section 4.2 and  Appendix D), except for $\alpha$. Results for $\alpha=\{0,0.25,0.5,0.75\}$ are enumerated in Table S11. The last row contains the performance of the seasonal persistence model (denoted as “Naïve”) with a seasonality set to $168$ time steps ($1$ week).
@@
E.5 Computational cost of Sonnet
@@ Appendix F Limitations
Appendix G Social Impact
Although it is difficult to predict the full scope of Sonnet’s influence, we can anticipate both beneficial and potentially adverse social implications arising from its adoption. On the positive side, enhanced time series forecasting—especially when grounded in realistic, application-oriented evaluation protocols—can substantially improve decision-making across a wide array of domains, including public health, energy demand and finance. By providing reliable projections, our approach may lead to more efficient use of scarce resources, better anticipation of system stresses and, ultimately, improved societal welfare. Furthermore, rigorous benchmarking against real-world scenarios helps to bridge the gap between academic research and practical deployment, encouraging translational collaborations that can accelerate the uptake of forecasting innovations in industry and government.
Conversely, over-reliance on algorithmic forecasts carries inherent risks. No matter how sophisticated, forecasting models (including Sonnet) cannot fully account for unforeseen systemic shocks, data biases or shifts in underlying processes. If taken as infallible oracles, such predictions may engender misplaced confidence, leading to suboptimal or even harmful policy and operational choices. To mitigate these dangers, it is essential that model outputs be interpreted judiciously by domain experts, who can contextualise quantitative predictions within broader qualitative insights.
Appendix H Supplementary ILI rate forecasting figures

### v2

@@ E.2 Supplementary experiment results with sMAPE score
$\mu$ 1.60 1.46 2.92 2.01 2.00
$\sigma$ 0.07 0.14 0.07 0.20 0.07
@@
$\mu$ 2.23 1.69 4.40 2.88 2.80
$\sigma$ 0.15 0.08 0.15 0.10 0.06
Table S11: Seed control (55 seeds) for Sonnet using the ILI-ENG data set across all forecasting horizons ($H$). $\mu$ and $\sigma$ denote the mean and standard deviation of the $5$ obtained MAEs per test period. The results in the main paper were obtained for seed ‘42’.
E.3 Ablation study on Sonnet
An ablation study is performed to better understand the contribution of Sonnet’s components to the overall performance. We ablate over the following modules: the coherence mechanism (within MVCA), the MLP head for inter-variable dependency modelling (within MVCA), the complete MVCA module (section 3.3), the joint embedding module (section 3.1), and the Koopman evolvement module (section 3.4).
@@ E.3 Ablation study on Sonnet
Table S12: Effect of different $\alpha$ values with Sonnet on the ELEC data set. $r$ and $\epsilon\%$ denote the correlation and sMAPE scores, respectively. Best results are bolded, second best underlined.
A detailed description of each ablation variant is provided as follows:
  • $\neg$ MLP: This denotes that we do not add the residual connection with the two-layer perceptron to obtain the attention output. Instead of computing $\mathbf{O}_{m}=\mathbf{O}_{r}+\mathrm{MLP}\left(\mathbf{O}_{r}\right)$, we directly have $\mathbf{O}_{m}=\mathbf{O}_{r}$. In this case, the inter-variable dependency via the MLP is omitted. Although some inter-variable interactions are partially (no learnable parameter is included in this process) captured when transforming the input into the frequency space along the feature dimension.
We re-tune Sonnet’s hyperparameters for each ablation variant across all locations, test seasons, and forecasting horizons (for the ILI and WEA tasks).
Results are enumerated in Table S10. We report the average MAE across all test seasons and its percentage of difference compared to the Sonnet model (enumerated in Table 1). Among all the ablated components, the MVCA module has the greatest impact. Removing MVCA results in an increase of MAE by $6.3\%$ on average across the ILI and WEA tasks. Breaking down MVCA further, we observe that removing its two internal components: coherency modelling and the MLP residual connection, leads to notable drops in performance ($5.2\%$ and $4.1\%$ increase in MAE, respectively), but less pronounced than removing the entire MVCA block. This indicates that while capturing coherence over time and enhancing variable interactions with residual connected MLP are both important modules, they also work in tandem within the full MVCA design.
E.4 Sensitivity to random seed initialisation
Results that assess the sensitivity of Sonnet to different random seed initialisations are enumerated in Table S11. We have examined $5$ different seeds (‘10’, ‘42’, ‘111’, ‘1111’, ‘1234’) while conducting experiments on the ILI-ENG data set for all test periods and forecasting horizons. The obtained results highlight that the proposed model provides consistent results across all regions and forecasting horizons. Results do not deviate significantly from the ones presented in the main paper (Table 1) which are obtained using seed ‘42’. Standard deviations in the averaged MAEs (all test periods) over the $5$ seeds for each forecasting horizon do not exceed $0.07$ which in real terms (using the units of the actual data) represents $7$ people in a population of $10$ million. This highlights the robustness of the mean estimate.
E.5 Analysis of model performance for tasks with strong seasonality
Task $H$ Sonnet DeformTime ModernTCN Samformer TimeXer PatchTST Itransformer Crossformer DLinear
ELEC
12 0.0885 0.1033 0.1440 0.1801 0.1114 0.1267 0.1261 0.1330 0.2630
24 0.1034 0.1223 0.1649 0.2198 0.1423 0.1368 0.1394 0.1703 0.2997
36 0.1138 0.1507 0.1788 0.2441 0.1546 0.1431 0.1554 0.1969 0.3259
ENER
24 0.3152 0.3474 0.3133 0.3256 0.3178 0.3262 0.3220 0.3205 0.3691
48 0.3588 0.3962 0.3692 0.3715 0.4186 0.3791 0.3808 0.4253 0.4246
72 0.3754 0.4071 0.3792 0.3814 0.4227 0.3867 0.3973 0.4123 0.4458
168 0.3796 0.4620 0.4175 0.3986 0.4196 0.4352 0.3991 0.4772 0.4604
ETTh1
96 0.1972 0.1840 0.1774 0.1905 0.1807 0.1757 0.1825 0.1989 0.2045
192 0.2187 0.1988 0.2013 0.2203 0.2043 0.1999 0.2051 0.2149 0.2834
336 0.2462 0.2024 0.2123 0.2433 0.2226 0.2248 0.2322 0.2577 0.3877
720 0.2661 0.2390 0.2288 0.2648 0.2292 0.2500 0.2449 0.3000 0.5007
ETTh2
96 0.2743 0.2924 0.2790 0.2986 0.2797 0.2765 0.2991 0.3254 0.2840
192 0.3185 0.3010 0.3183 0.3254 0.3331 0.3174 0.3463 0.3534 0.3313
336 0.3463 0.3172 0.3350 0.3622 0.3769 0.3417 0.3857 0.3891 0.4056
720 0.3729 0.3577 0.4259 0.4239 0.4069 0.3887 0.3945 0.4248 0.5391
ILI-ENG
7 1.3204 1.2802 1.7351 1.7701 1.7086 1.9402 1.7746 1.3243 2.1892
14 1.5964 1.6521 2.2656 2.1201 1.9703 2.4857 2.3660 1.9326 2.7452
21 2.1226 2.3246 2.3331 2.8860 2.8852 2.8458 3.0904 2.8733 3.2983
28 2.4956 2.7829 2.5948 3.3530 2.7894 3.2161 3.4784 3.0740 3.6805
ILI-US2
7 0.3795 0.4142 0.4523 0.5140 0.4931 0.6400 0.5455 0.4444 0.7214
14 0.4186 0.4544 0.5501 0.5784 0.5402 0.7302 0.6357 0.4695 0.7899
21 0.4914 0.5374 0.5576 0.6004 0.6190 0.7489 0.6554 0.5720 0.8642
28 0.5740 0.5554 0.5892 0.6561 0.6808 0.8450 0.7370 0.5842 0.8920
ILI-US9
7 0.2727 0.2614 0.2945 0.3396 0.3435 0.3882 0.3725 0.3222 0.4348
14 0.2925 0.3040 0.3398 0.3965 0.3715 0.4334 0.4049 0.3254 0.4780
21 0.2955 0.3187 0.3558 0.3811 0.4430 0.4664 0.4504 0.3285 0.5315
28 0.3671 0.3913 0.3786 0.4202 0.4774 0.5002 0.5535 0.4018 0.5622
WEA-CT
4 1.1748 1.3217 1.3716 1.6459 1.9038 2.8087 1.5908 1.1470 2.2270
12 2.4586 2.5082 2.6260 2.9155 2.8796 3.6686 2.8671 2.3716 3.3320
28 3.2135 3.3330 3.4816 3.4502 3.4469 3.6743 3.4602 3.2349 3.6682
120 3.6790 3.7547 3.9605 3.8332 3.7892 3.9108 3.8934 3.6567 3.9094
WEA-HK
4 0.5456 0.6001 0.5729 0.6577 0.7984 0.9391 0.6449 0.5435 0.7752
12 0.8753 0.9935 0.9314 0.9964 1.0159 1.2814 1.0259 0.8953 1.1804
28 1.1809 1.2712 1.2761 1.3163 1.2920 1.3973 1.3405 1.2660 1.3926
120 1.5093 1.6400 1.5321 1.6370 1.6032 1.7110 1.5684 1.5196 1.6268
WEA-LD
4 1.2448 1.4040 1.4175 1.5651 1.9055 2.2429 1.5485 1.2493 1.8231
12 2.2391 2.3942 2.3661 2.4950 2.5732 2.9561 2.5287 2.2368 2.7131
28 2.8627 3.0055 2.9744 3.0418 3.0354 3.2615 3.1209 2.8913 3.1500
120 3.2539 3.3166 3.6610 3.5329 3.5188 3.7507 3.6614 3.3583 3.6064
WEA-NY
4 0.9584 1.1266 1.0886 1.2500 1.5107 1.7682 1.2372 0.9673 1.4818
12 1.7458 1.8459 1.8486 1.9756 2.1140 2.3498 1.9589 1.7531 2.2809
28 2.2952 2.4040 2.4314 2.5138 2.5193 2.6537 2.5108 2.3223 2.7166
120 2.6635 2.8076 2.9735 2.9731 2.9984 2.9692 2.7759 2.7372 3.1808
WEA-SG
4 0.3109 0.3229 0.3243 0.3622 0.3596 0.4010 0.3542 0.3132 0.3809
12 0.3664 0.3849 0.3899 0.4189 0.4020 0.4487 0.4099 0.3765 0.4285
28 0.4213 0.4540 0.4565 0.4713 0.4475 0.4866 0.4655 0.4414 0.4719
120 0.4747 0.5128 0.5101 0.5018 0.5023 0.5191 0.4936 0.4980 0.5151
Table S13: Mean Absolute Error (MAE) evaluated over the full predicted sequence for all forecasting horizons ($H$), tasks, and models. In consistency with the main experimental setup, error metrics are averaged over 44 test seasons for ILI tasks, 33 years for the WEA tasks, and $2$ for ELEC. Best results are bolded, second best underlined.
@@ E.5 Analysis of model performance for tasks with strong seasonality
Prior work has shown that modelling inter-variable dependencies in time series problems may occasionally cause overfitting (nie2023time). While our results for the ILI and WEA forecasting tasks do not corroborate this (Section 4.2), we conduct further experiments on a different task (ELEC) that displays a different pattern of performance across most models. In these experiments, we explore different values for the parameter $\alpha$ in Sonnet. $\alpha$ controls the relative contribution of exogenous variables in the learned representations (section 3.1). For $\alpha=0$, Sonnet only uses historical values of the target (endogenous) variable. We use the ELEC data set for evaluation because for this specific task, PatchTST—that does not capture inter-variable dependencies—outperforms some models that do (e.g. ModernTCN and Crossformer). In addition, $7$ out of $9$ evaluated models cannot surpass the accuracy of the seasonal persistence model (both in terms of MAE and sMAPE), indicating that for ELEC, seasonal patterns are consistent over time. Sonnet’s hyperparameters are re-tuned (see Section 4.2 and Appendix D), except for $\alpha$. Results for $\alpha=\{0,0.25,0.5,0.75\}$ are enumerated in  S12. The last row contains the performance of the seasonal persistence model (denoted as “Naïve”) with a seasonality set to $168$ time steps ($1$ week).
@@
E.6 Computational cost of Sonnet
@@
E.7 Note on performance evaluation
We acknowledge that most prior work in time series forecasting evaluates model performance by averaging errors across the entire prediction horizon. While this provides a general view of forecasting accuracy, it may obscure performance at the target time step. As the prediction horizon extends, forecasting becomes more challenging (jhyndman2006another). Model performance can thus vary significantly between forecasting the closer / further time steps. For example, consider two models whose predictions over a three-step sequence yield MAEs of $\{0.1,0.2,0.6\}$ and $\{0.2,0.3,0.4\}$, respectively. While both models have the same averaged MAE when evaluated over the sequence, the second model performs better on the final time step, which is the main forecasting target. Averaging errors over all time steps may benefit models that excel in early predictions but struggle near the target horizon. We further demonstrate this point in the coming section with empirical results.
E.8 Evaluating forecasts over the entire output sequence
To provide a more comprehensive understanding of model performance, we also provide the forecasting accuracy by averaging the error across the entire output sequence, as opposed to just using the target time step (Table 1). The results are presented in Table S13. It can be observed that Sonnet ranks first in $31$ over $47$ tasks and second in $9$ tasks. The averaged MAE reduction is $1.0\%$. Additionally, Sonnet reduces the MAE of DeformTime, which is the best-performing baseline, by $4.7\%$.
For the ILI and WEA data sets, the best-performing baseline is DeformTime when evaluated at the target time step, with Sonnet reducing the MAE by $3.3\%$ on average. However, Crossformer is the best performing model when evaluated over the sequence, with Sonnet reducing the MAE by $5.4\%$. This observation indicates that DeformTime performs better than Crossformer at the target horizon, while Crossformer yields better performance when considering the entire series of forecasts. This change in the ranking of baselines supports our argument in Section E.7 that averaging errors across the sequence, rather than evaluating the error at the target forecasting horizon, can lead to a misleading assessment of how the model performs at different forecasting horizons.
@@ Appendix F Limitations
Appendix G Supplementary ILI rate forecasting figures